A Slinky has a k-value of 0.80 kg/s2.  If it is stretched to 2.7 meters, what is its elastic potential energy?


1
Expert's answer
2021-02-05T02:13:47-0500
"PE_{elastic}=\dfrac{1}{2}kx^2,""PE_{elastic}=\dfrac{1}{2}\cdot 0.8\ \dfrac{kg}{s^2}\cdot(2.7\ m)^2=2.92\ J."